OpenStudy (anonymous):

Solve multistory equations 2/3x + 1/2 = 4/3x - 1/6

OpenStudy (anonymous):

Multistep

OpenStudy (campbell_st):

add 1/6 to both sides 2/3 x + 2/3 = 4/3 x then subtract 2/3 x from both sides.. 2/3 = 2/3x this should make it easier to solve